# TRA: Better Length Generalisation with Threshold Relative Attention

### v2

@@ Abstract
Transformers struggle with length generalisation, displaying poor performance even on basic tasks. We test whether these limitations can be explained through two key failures of the self-attention mechanism. The first is the inability to fully remove irrelevant information. The second is tied to position, even if the dot product between a key and query is highly negative (i.e. an irrelevant key) learned positional biases may unintentionally up-weight such information - dangerous when distances become out of distribution. Put together, these two failure cases lead to compounding generalisation difficulties. We test whether they can be mitigated through the combination of a) selective sparsity - completely removing irrelevant keys from the attention softmax and b) contextualised relative distance - distance is only considered as between the query and the keys that matter. We show how refactoring the attention mechanism with these two mitigations in place can substantially improve generalisation capabilities of decoder only transformers. [^1] Code available at: https://github.com/mopper97/TRA
@@ 2.4 The Final Attention Weights
This gives us the full TRA mechanism. Note keys which do not meet the threshold are completely masked out from the final attention softmax, and distances are only considered between the query and the keys that matter. This means that TRA exhibits both selective sparsity and allows semantic content to determine and consequently synergise with the positional bias.
@@ 3.2 Baselines and Experimental Setup
Forgetting Transformer (FoT): The recent Forgetting Transformer (FoT) (Lin et al., 2025) utilises a forget gate as its positional encoding . The crucial difference between it and TRA is that the forgetting transformer uses standard relative distance compared with TRA contextualised alternative. It therefore serves as the truest test as to whether the issues we hypothesise regarding position are valid.
@@ 3.4 Flip-Flops++
Results are shown in Table 2. As before most PE methods struggle to generalise, though RoPE performs quite strongly with first match, echoing the findings of Barbero et al. (2025). However, in this more challenging setting, only the contextualised schemes, CoPE and TRA, are able to robustly generalise. Though TRA performs better at last match while CoPE holds a slight edge at first match, due to their use of different positional schemes. For natural language it is likely that the former is preferable as it better reflects causal structure (i.e. it allows understanding of an entities current state). Notably FoT, which in the original Flip-Flop benchmark of Liu et al. (2023a) appeared to be a strong contender, suffers substantial performance degradation in this more challenging setting. This is because, as we hypothesised in the introduction, it is unable to synergise positional and semantic information and consequently fails to generalise to capability (3) from earlier. This shortcoming is clearly visible by contrasting the attention weights of FoT and TRA.

### v3

@@ Abstract
Transformers struggle with length generalisation, displaying poor performance even on basic tasks. We test whether these limitations can be explained through two key failures of the self-attention mechanism. The first is the inability to fully remove irrelevant information. The second is tied to position, even if the dot product between a key and query is highly negative (i.e. an irrelevant key) learned positional biases may unintentionally up-weight such information - dangerous when distances become out of distribution. Put together, these two failure cases lead to compounding generalisation difficulties. We test whether they can be mitigated through the combination of a) selective sparsity - completely removing irrelevant keys from the attention softmax and b) contextualised relative distance - distance is only considered as between the query and the keys that matter. We show how refactoring the attention mechanism with these two mitigations in place can substantially improve generalisation capabilities of decoder only transformers. [^1] Full code available at: https://github.com/microsoft/tra. Core mechanism also in Appendix B.
@@ 2.4 The Final Attention Weights
This gives us the full TRA mechanism. We use $\log(\bm{D^{\prime}})$ for improved numerical stability and weight scale following Lin et al. (2025). Note keys which do not meet the threshold are completely removed from the final softmax and do not count towards distance. This means that TRA exhibits both selective sparsity and allows semantic content to determine, and consequently synergise with, positional bias.
@@ 3.2 Baselines and Experimental Setup
Forgetting Transformer (FoT): The recent Forgetting Transformer (FoT) (Lin et al., 2025) utilises a forget gate as its positional encoding. The crucial difference between it and TRA is that the forgetting transformer uses standard relative distance compared with TRA’s contextualised alternative. It therefore serves as the truest test as to whether the issues we hypothesise regarding position are valid.
@@ 3.4 Flip-Flops++
Results are shown in Table 2. As before most PE methods struggle to generalise, though RoPE performs quite strongly with first match, echoing the findings of Barbero et al. (2025). However, in this more challenging setting, only the contextualised schemes, CoPE and TRA, are able to robustly generalise. Though TRA performs better at last match while CoPE holds a slight edge at first match, due to their use of different positional schemes. For natural language it is likely that the former is preferable as it better reflects causal structure (i.e. it allows understanding of an entity’s current state). Notably FoT, which in the original Flip-Flop benchmark of Liu et al. (2023a) appeared to be a strong contender, suffers substantial performance degradation in this more challenging setting. This is because, as we hypothesised in the introduction, it is unable to synergise positional and semantic information and consequently fails to generalise to capability (3) from earlier. This shortcoming is clearly visible by contrasting the attention weights of FoT and TRA.
@@ Appendix A Hyperparameters:
Appendix B Implementation:
1class TRA(nn.Module):
2 def __init__(self, e_dim, n_heads, dropout=0.01):
3 super(TRA, self).__init__()
4 self.E = e_dim
5 self.e = e_dim // n_heads
6 self.h = n_heads
7 self.tokeys = nn.Linear(e_dim, e_dim, bias=False)
8 self.toqueries = nn.Linear(e_dim, e_dim, bias=False)
9 self.tovalues = nn.Linear(e_dim, e_dim, bias=False)
10 self.delta_proj = nn.Linear(e_dim, self.h)
11 self.c_out = nn.Linear(e_dim, e_dim)
12 # use qk norm but no scale to save params
13 self.qk_norm = RMSNorm(self.e, elementwise_affine=False)
14 self.register_buffer("bias", torch.triu(torch.ones(max_len, max_len), diagonal=1).view(1, 1,max_len, max_len).to(torch.bool))
15 self.dropout = nn.Dropout(dropout)
16
17 def forward(self, x):
18 b, s, _ = x.shape # batch_size, seq_len, e_dim
19 # transform q, k, v
20 q = self.toqueries(x).view(b, s, self.h, self.e).transpose(1, 2)
21 k = self.tokeys(x).view(b, s, self.h, self.e).transpose(1, 2)
22 v = self.tovalues(x).view(b, s, self.h, self.e).transpose(1, 2)
23 # qk norm (no scale -> can share)
24 q = self.qk_norm(q)
25 k = self.qk_norm(k)
26 # compute attn dot
27 S = (q @ k.transpose(-2,-1)) / torch.sqrt(torch.tensor(self.e, dtype=torch.float, device=x.device))
28 S = S.masked_fill(self.bias[:,:,:s,:s], float(’-1e11’))
29 S = F.relu(S)
30 mask = (S == 0)
31 # compute positional weights D
32 delta = F.logsigmoid(self.delta_proj(x)).transpose(1,2).unsqueeze(-1).contiguous().view(b, self.h, s, 1)
33 D = (~mask).float() * delta
34 D = D.flip(-1).cumsum(-1).flip(-1)
35 # dropout crucial to avoid loss spikes in language modelling
36 A = self.dropout(S + D)
37 A[mask] = -1e11 # mask: casual + threshold
38 A = A.softmax(dim=-1)
39 # protect against leaks where mask all = 0 from no-op attention
40 A = A.masked_fill(mask.all(-1).unsqueeze(-1), 0)
41 # compute output
42 out = A @ v
43 out = out.transpose(1, 2).contiguous().view(b, s, self.E)
44 return self.c_out(out)
Listing 1: TRA Implementation